What is Lorentz number? The thermal and electrical conductivities of Cu at 20oC are 390 Wm-1K-1 and

5.87x107W-1m-1 respectively. Calculate Lorentz number.​​


1
Expert's answer
2022-05-02T13:44:15-0400

The Lorentz number is given by

"L=\frac{\kappa}{\sigma T}""L=\rm\frac{290\: Wm^{-1}K^{-1}}{5.87\times 10^{7}\: \Omega ^{-1}m^{-1}\times 293\: K}=1.69\times 10^{-8}\: W\Omega K^{-2}"